# Quantum framework for Reinforcement Learning: integrating Markov Decision Process, quantum arithmetic, and trajectory search

### v2

@@ Abstract
This paper introduces a quantum framework for addressing reinforcement learning (RL) tasks, grounded in the quantum principles and leveraging a fully quantum model of the classical Markov Decision Process (MDP). By employing quantum concepts and a quantum search algorithm, this work presents the implementation and optimization of the agent-environment interactions entirely within the quantum domain, eliminating reliance on classical computations. Key contributions include the quantum-based state transitions, return calculation, and trajectory search mechanism that utilize quantum principles to demonstrate the realization of RL processes through quantum phenomena. The implementation emphasizes the fundamental role of quantum superposition in enhancing computational efficiency for RL tasks. Results demonstrate the capacity of a quantum model to achieve quantum enhancement in RL, highlighting the potential of fully quantum implementations in decision-making tasks. This work not only underscores the applicability of quantum computing in machine learning but also contributes the field of quantum reinforcement learning (QRL) by offering a robust framework for understanding and exploiting quantum computing in RL systems.
@@ II.4 Full quantum reinforcement learning
To further advance fully quantum RL, our framework enables direct interaction between the quantum agent and the environment entirely within the quantum domain. The concept of a quantum-accessible environment was first introduced in [36] where the agent and environment interact through an intermediate register. In comparison, our method enables direct communication between the agent and environment, eliminating the need for an additional layer. This direct interaction allows for the simultaneous evaluation of multiple states and actions, increasing computational parallelism and enabling faster exploration of the state-action space. Furthermore, since these interactions are entirely within the quantum domain, this will remove the classical-quantum conversions and achieve the complete quantum interactions. Our framework facilitates optimization by a quantum search algorithm directly to the quantum samples generated through these interactions. This approach eliminates reliance on classical subroutines and achieve significant computational speedups. To the best of our knowledge, our work is the first attempt to implement a complete quantum framework for reinforcement learning by integrating the quantum realization of the classical Markov Decision Process (MDP).
@@ IV.1.1 Agent and environment on quantum computer
We are focusing on a classical MDP circuit comprising four states $S=\{s_{0},s_{1},s_{2},s_{3}\}$ and two actions $A=\{a_{0},a_{1}\}$. In the quantum implementation of this MDP, we define a state space S and an action space A. For the state space S, each state $s\in S$ is encoded as an orthonormal basis vector in a Hilbert space and expressed as a quantum state |$\ket{s}$. Similarly, all the possible actions are represented as vectors |$\ket{a}\in A$ in the action space A. To represent N distinct states in a quantum system, n qubits are required, where $n=\log_{2}(N)$, to encode all the possible states into quantum samples. Therefore, our quantum MDP necessitates two qubits for the state set S and one qubit for the action set A. These qubits can be described as strings of qubits on a quantum computer, where each string represents the binary representation of an integer corresponding to that quantum sample, i.e.,
@@
As we are implementing MDP in the quantum domain, these states and actions can be prepared into superposition by using Hadamard transformation. This process can initialize qubits as a uniform combination of quantum samples for all states and actions within their respective spaces. For the state space S, we define the initial ground state of each state register as |$\ket{0_{s}}=\ket{0}$. Applying the Hadamard transformation to these registers generates a uniform superposition of quantum states, as formulated in Eq. (2).
@@
After preparing the state S and action A into uniform superposition quantum samples, we then consider the state transition function, which defines the likelihood of transitioning from one state to another given a specific action. In a classical MDP, the state transition function $P(s^{\prime}|s,a)$ determines the probability of transitioning to next state ${s^{\prime}}$ given the current state $s$ and action $a$. In a quantum circuit, these probabilities are encoded in the amplitudes of quantum states, where the squared magnitudes of the amplitudes correspond to probabilities. This is achieved by applying $R_{y}(\theta)$ to an ancillary qubit initialized in the |$\ket{0}$ state, producing a superposition in which the probabilities of measuring |$\ket{0}$ and |$\ket{1}$ are $\cos^{2}\left(\frac{\theta}{2}\right)\text{ and }\sin^{2}\left(\frac{\theta}{2}\right)$, respectively. To encode classical transition probabilities $P(s^{\prime}|s,a)$, the rotation angle $\theta$ is calculated as $\theta=2\arcsin(\sqrt{P(s^{\prime}|s,a)})$. In our work, we use multi-controlled $R_{y}(\theta)$ gates to apply these rotations conditionally based on the current state |$\ket{s}$ and action |$\ket{a}$. The rotation is applied only when the control qubits match a specific state-action pair. Equation (6) is a mathematically representation of this quantum state transition process where the controlled $R_{y}(\theta)$ gates apply a rotation to an ancillary qubit |$\ket{0_{s^{\prime}}}$ by an angle $\theta$, depending on the state-action pair, thereby determining the next state, |$\ket{s^{\prime}}\in S$.
@@
The reward function in a classical MDP is defined as $R(s,a)$, where specific state-action pairs $(s,a)$ determine a reward value. The Controlled-NOT (CNOT) gate models this behavior conditionally, where the control qubits (representing the current state |$\ket{s}$ and action |$\ket{a}$) dictate whether the target qubit (representing the reward |$\ket{r}$ ) is flipped. In our quantum circuit, the next state qubits |$\ket{s^{\prime}}$ are used as control qubits since they represent the result of transitioning from a specific state-action pair to a next state |$\ket{s^{\prime}}$, implicitly encoding the state-action pair $(s,a)$ in determining the reward. The CNOT gate works conditionally such that the reward qubit |$\ket{r}$, initially set to |$\ket{0}$, is flipped to |$\ket{1}$ if the next state matches a condition where a reward exists. This ensures that the reward qubit reflects the reward function based on the transition outcome (next state), determined by the state-action pair. The mathematical formulation of the reward function can be expressed as Eq. (7) where the CNOT gate flips an ancillary qubit |$\ket{0_{r}}$ when the control qubits match valid next states |$\ket{s^{\prime}}$, thereby generating the corresponding reward quantum sample.
@@ IV.1.2 Agent and environment single interaction on quantum computer
At each interaction step, the agent receives the current quantum state |$\ket{s}$ of the environment. Then, the agent applies a unitary operation that corresponds to its policy $\pi$ encoded as |$\ket{a}$. This process of action selection influences the evolution of the environment’s state, resulting in a transition of the system to the next quantum state |$\ket{s^{\prime}}$. Then, the agent receives the reward |$\ket{r}$ based on the next state |$\ket{s^{\prime}}$. In the context of QMDP, this agent-environment interaction can be modeled using a unitary operator $U(S\otimes A\otimes S\otimes R)$, which prepares a quantum state representing the distribution of trajectories for the state, action, next state and reward sequence. This quantum state can be expressed as follows in Eq. (8).
@@ IV.2 Agent-environment quantum interactions over multiple time steps
To extend the QMDP framework over T time steps, we need to account for the quantum state transition and the calculation of cumulative rewards. At each time step t, the agent receives a quantum state |$\ket{s_{t}}$, performs an action |$\ket{a_{t}}$, and the environment responds by transitioning to a new state |$\ket{s^{\prime}_{t}}$ while providing a reward |$\ket{r_{t}}$. This process is repeated over T time steps, ensuring that the next state |$\ket{s^{\prime}_{t}}$ at t becomes the current state |$\ket{s_{t}}$ for the subsequent time step.
@@
For the $T$ time steps of interaction, each quantum sample can be represented as a string of qubits encoding the possible transitions from ${t=0}$ to ${t=T-1}$, formed by concatenating state-action-next state-reward sequences. This is obtained by applying a sequence of individual interactions. The complete quantum distribution of the agent-environment interaction at time step $T$ can be summarized as Eq. (10).
@@ IV.2.1 Return calculation by quantum arithmetic
In our work, this classical concept is mapped into the quantum domain, where the return is encoded as a quantum state |$\ket{g}$ in a Hilbert space $G$. This space represents a quantum system with a sufficient number of qubits to encode every potential return value. Initially, the quantum state |$\ket{g}$ is prepared in its ground state |$\ket{0_{g}}$. To compute return, the quantum operation $U_{G}$ is applied to the reward qubits |$\ket{r_{t}}$ for each time step. This operation executes a quantum addition that performs sequential bit-wise addition of each reward register |$\ket{r_{t}}$ and stores the result in the return registers |$\ket{g}$. This operation propagates carry bits through the circuit, similar to a classical addition. For each reward register |$\ket{r_{t}}$, a series of controlled-NOT and Toffoli gates are applied to implement bit-wise addition into the return registers. Conceptually, this operation accumulates the discounted sum of rewards across T time steps. An abstract representation of this transformation is given in Eq. (11), where $\gamma\in[0,1]$ is a fixed discount factor, $T$ is the total number of time steps, and |$\ket{r_{t}}$ represents the reward at time step $t$.
@@ IV.3 Quantum trajectories search
In quantum reinforcement learning (QRL), we apply the power of quantum algorithms to efficiently search for desired trajectories and identify the best action for each state. We utilize Grover’s search [12] to identify optimal quantum trajectories from a set of agent-environment interactions over multiple time steps. Each trajectory consists of the agent’s action, the resulting quantum state of the environment, and a return quantum sample. The algorithm seeks to find optimal quantum trajectories, which can inform the optimal policies for a quantum reinforcement learning (QRL) agent.
We formulated the interaction between a quantum agent and its environment over $T$ time steps (Eq. (12)). A quantum trajectory here refers to the series of quantum states and actions the agent takes, together with the environment’s responses. Each trajectory |$\ket{T}$ can be represented as |$\ket{T}=\ket{s_{0},a_{0},s^{\prime}_{0},r_{0},s_{1},a_{1},s^{\prime}_{1},r_{1},\ldots,s_{T-1},a_{T-1},s^{\prime}_{T-1},r_{T-1},g}$ where $s$ is the current state at each time step from $t=0$ to $t=T-1$, $a$ is the action taken by the agent, $s^{\prime}$ is the next state reached after taking an action at state $s$, $r$ is the reward received and $g$ is the total return throughout the trajectory.
@@ V Demonstrations and Results
In the demonstrations and results section of this work, we first present the dynamics of the classical Markov Decision Process (MDP) through a graphical representation of state transitions and associated probabilities, as shown in Fig. 3. The diagram includes four states labeled $s_{0},s_{1},s_{2}$ and $s_{3}$, and two actions, $a_{0}$ and $a_{1}$. The arrows connecting the states indicate the transitions between these states for a given action. Each transition is labeled with a probability value that denotes the likelihood of moving from one state to another under a specific action. These probabilities provide insights into the stochastic nature of the state transitions. For example, transitioning from $s_{0}$ to $s_{1}$ with action $a_{0}$ has a probability of 0.6. Similarly, transitioning from $s_{0}$ to $s_{2}$ under action $a_{0}$ occurs with a probability of 0.4. The diagram also illustrates several self-transitions, such as the one in state $s_{3}$, where action $a_{1}$ leads to a transition back to $s_{3}$ with a probability of 1. These self-transitions capture situations where the system remains in the same state regardless of the action taken. The figure visualizes the probabilistic behavior of the system under different actions, as well as how the state evolves over time based on these actions.
@@ V.1.1 Agent and environment on quantum computer
Figure 4 illustrates the Quantum Markov Decision Process (QMDP), where the agent’s actions and the environment’s responses are represented by the quantum state.
@@ V.1.2 Agent and environment single interaction on quantum computer
To verify the accuracy of the Quantum Markov Decision Process (QMDP) circuit, we simulate a single interaction between the agent and the environment on a quantum computer. Using both a state transition heat-map and a quantum sample distribution plot, we demonstrate that the QMDP circuit correctly mirrors the state transitions and rewards of the classical MDP, mentioned in Fig. 3.
@@
Figure 6: Quantum sample distribution of the QMDP circuit, displaying the probability of measuring each quantum state after simulating a single agent-environment interaction. The x-axis lists possible quantum states, while the y-axis indicates their probabilities.
Figure 6 illustrates the quantum sample distribution of the QMDP circuit, displaying the probability of observing each possible quantum sample after a single execution of the circuit. Each quantum sample in binary notation (on the x-axis) represents a combination of the agent’s initial state, chosen action, resulting next state and reward sequence encoded in different qubits. Table 1 presents the qubit representations of each quantum sample. The y-axis indicates the probability of each quantum sample. Peaks in this distribution, quantum samples |$\ket{1111111}$ and |$\ket{0101100}$ with probabilities 0.354 and 0.335 respectively, highlight high-probability sequences of this interaction. These results validate the QMDP circuit’s accurate representation of both state transition probabilities and reward during a single interaction.

### v3

@@ Abstract
This paper introduces a quantum framework for addressing reinforcement learning (RL) tasks, grounded in the quantum principles and leveraging a fully quantum model of the classical Markov decision process (MDP). By employing quantum concepts and a quantum search algorithm, this work presents the implementation and optimization of the agent-environment interactions entirely within the quantum domain, eliminating reliance on classical computations. Key contributions include the quantum-based state transitions, return calculation, and trajectory search mechanism that utilize quantum principles to demonstrate the realization of RL processes through quantum phenomena. The implementation emphasizes the fundamental role of quantum superposition in enhancing computational efficiency for RL tasks. Results demonstrate the capacity of a quantum model to achieve quantum enhancement in RL, highlighting the potential of fully quantum implementations in decision-making tasks. This work not only underscores the applicability of quantum computing in machine learning but also contributes to the field of quantum reinforcement learning (QRL) by offering a robust framework for understanding and exploiting quantum computing in RL systems.
@@ II.4 Full quantum reinforcement learning
To further advance fully quantum RL, our framework enables direct interaction between the quantum agent and the environment entirely within the quantum domain. The concept of a quantum-accessible environment was first introduced in [36] where the agent and environment interact through an intermediate register. In comparison, our method enables direct communication between the agent and environment, eliminating the need for an additional layer. This direct interaction allows for the simultaneous evaluation of multiple states and actions, increasing computational parallelism and enabling faster exploration of the state-action space. Furthermore, since these interactions are entirely within the quantum domain, this will remove the classical-quantum conversions and achieve the complete quantum interactions. Our framework facilitates optimization by a quantum search algorithm directly to the quantum states generated through these interactions. This approach eliminates reliance on classical subroutines and achieve significant computational speedups. To the best of our knowledge, our work is the first attempt to implement a complete quantum framework for reinforcement learning by integrating the quantum realization of the classical Markov decision process (MDP).
@@ IV.1.1 Agent and environment on quantum computer
We are focusing on a classical MDP circuit comprising four states $S=\{s_{0},s_{1},s_{2},s_{3}\}$ and two actions $A=\{a_{0},a_{1}\}$. In the quantum implementation of this MDP, we define a state space S and an action space A. For the state space S, each state $s\in S$ is encoded as an orthonormal basis vector in a Hilbert space and expressed as a quantum state |$\ket{s}$. Similarly, all the possible actions are represented as vectors |$\ket{a}\in A$ in the action space A. To represent N distinct states in a quantum system, n qubits are required, where $n=\log_{2}(N)$, to encode all the possible states into quantum states. Therefore, our quantum MDP necessitates two qubits for the state set S and one qubit for the action set A. These qubits can be described as strings of qubits on a quantum computer, where each string represents the binary representation of an integer corresponding to that quantum state, i.e.,
@@
As we are implementing MDP in the quantum domain, these states and actions can be prepared into superposition by using Hadamard transformation. This process can initialize qubits as a uniform combination of quantum states for all states and actions within their respective spaces. For the state space S, we define the initial ground state of each state register as |$\ket{0_{s}}=\ket{0}$. Applying the Hadamard transformation to these registers generates a uniform superposition of quantum states, as formulated in Eq. (2).
@@
After preparing all states in the state space S and all actions in the action space A as uniform superposition quantum states, we then consider the state transition function, which defines the likelihood of transitioning from one state to another given a specific action. In a classical MDP, the state transition function $P(s^{\prime}|s,a)$ determines the probability of transitioning to next state ${s^{\prime}}$ given the current state $s$ and action $a$. In a quantum circuit, these probabilities are encoded in the amplitudes of quantum states, where the squared magnitudes of the amplitudes correspond to probabilities. This is achieved by applying $R_{y}(\theta)$ to an ancillary qubit initialized in the |$\ket{0}$ state, producing a superposition in which the probabilities of measuring |$\ket{0}$ and |$\ket{1}$ are $\cos^{2}\left(\frac{\theta}{2}\right)\text{ and }\sin^{2}\left(\frac{\theta}{2}\right)$, respectively. To encode classical transition probabilities $P(s^{\prime}|s,a)$, the rotation angle $\theta$ is calculated as $\theta=2\arcsin(\sqrt{P(s^{\prime}|s,a)})$. In our work, we use multi-controlled $R_{y}(\theta)$ gates to apply these rotations conditionally based on the current state |$\ket{s}$ and action |$\ket{a}$. The rotation is applied only when the control qubits match a specific state-action pair (|$(\ket{s^{*}},\ket{a^{*}})$, where |$\ket{s^{*}}\in\textit{S}$, |$\ket{a^{*}}\in\textit{A}$. Equation (6) is a mathematical representation of this quantum state transition process where the controlled $R_{y}(\theta)$ gates apply a rotation to an ancillary qubit |$\ket{0_{s^{\prime}}}$ by an angle $\theta$, depending on the state-action pair, thereby determining the next state, |$\ket{s^{\prime}}\in S$.
@@
The reward function in a classical MDP is defined as $R(s,a)$, where specific state-action pairs $(s,a)$ determine a reward value. The Controlled-NOT (CNOT) gate models this behavior conditionally, where the control qubits (representing the current state |$\ket{s}$ and action |$\ket{a}$) dictate whether the target qubit (representing the reward |$\ket{r}$ ) is flipped. In our quantum circuit, the next state qubits |$\ket{s^{\prime}}$ are used as control qubits since they represent the result of transitioning from a specific state-action pair to a next state |$\ket{s^{\prime}}$, implicitly encoding the state-action pair $(s,a)$ in determining the reward. The CNOT gate works conditionally such that the reward qubit |$\ket{r}$, initially set to |$\ket{0}$, is flipped to |$\ket{1}$ if the next state matches a condition where a reward exists. This ensures that the reward qubit reflects the reward function based on the transition outcome (next state), determined by the state-action pair. The mathematical formulation of the reward function can be expressed as Eq. (7) where the CNOT gate flips an ancillary qubit |$\ket{0_{r}}$ when the control qubits match valid next states |$\ket{s^{\prime}}$, thereby generating the corresponding reward.
@@ IV.1.2 Agent and environment single interaction on quantum computer
At each interaction step, the agent receives the current state |$\ket{s}$ of the environment. Then, the agent applies a unitary operation that corresponds to its policy $\pi$ encoded as |$\ket{a}$. This process of action selection influences the evolution of the environment’s state, resulting in a transition of the system to the next state |$\ket{s^{\prime}}$. Then, the agent receives the reward |$\ket{r}$ based on the next state |$\ket{s^{\prime}}$. In the context of QMDP, this agent-environment interaction can be modeled using a unitary operator $U(S\otimes A\otimes S\otimes R)$, which prepares a quantum state representing the distribution of trajectories for the state, action, next state and reward sequence. This quantum state can be expressed as follows in Eq. (8).
@@ IV.2 Agent-environment quantum interactions over multiple time steps
To extend the QMDP framework over T time steps, we need to account for the quantum state transition and the calculation of cumulative rewards. At each time step t, the agent receives a state |$\ket{s_{t}}$, performs an action |$\ket{a_{t}}$, and the environment responds by transitioning to a new state |$\ket{s^{\prime}_{t}}$ while providing a reward |$\ket{r_{t}}$. This process is repeated over T time steps, ensuring that the next state |$\ket{s^{\prime}_{t}}$ at t becomes the current state |$\ket{s_{t}}$ for the subsequent time step.
@@
For the $T$ time steps of interaction, each quantum state can be represented as a string of qubits encoding the possible transitions from ${t=0}$ to ${t=T-1}$, formed by concatenating state-action-next state-reward sequences. This is obtained by applying a sequence of individual interactions. The complete quantum distribution of the agent-environment interaction at time step $T$ can be summarized as Eq. (10).
@@ IV.2.1 Return calculation by quantum arithmetic
In our work, this classical concept is mapped into the quantum domain, where the return is encoded as a quantum state |$\ket{g}$ in a Hilbert space $G$. This space represents a quantum system with a sufficient number of qubits to encode every potential return value. Initially, the quantum state |$\ket{g}$ is prepared in its ground state |$\ket{0_{g}}$. To compute return, the quantum operation $U_{G}$ is applied to the reward qubits |$\ket{r_{t}}$ for each time step. This operation executes a quantum addition that performs sequential bit-wise addition of each reward register |$\ket{r_{t}}$ and stores the result in the return registers |$\ket{g}$. This operation propagates carry bits through the circuit, similar to a classical addition. For each reward register |$\ket{r_{t}}$, a series of CNOT and Toffoli gates are applied to implement bit-wise addition into the return registers. Conceptually, this operation accumulates the discounted sum of rewards across T time steps. An abstract representation of this transformation is given in Eq. (11), where $\gamma\in[0,1]$ is a fixed discount factor, $T$ is the total number of time steps, and |$\ket{r_{t}}$ represents the reward at time step $t$.
@@ IV.3 Quantum trajectories search
In quantum reinforcement learning (QRL), we apply the power of quantum algorithms to efficiently search for desired trajectories and identify the best action for each state. We utilize Grover’s search [12] to identify optimal quantum trajectories from a set of agent-environment interactions over multiple time steps. Each trajectory consists of the agent’s action, the resulting quantum state of the environment, and a return. The algorithm seeks to find optimal quantum trajectories, which can inform the optimal policies for a quantum reinforcement learning (QRL) agent.
We formulated the interaction between a quantum agent and its environment over $T$ time steps (Eq. (12)). A quantum trajectory here refers to the series of states and actions the agent takes, together with the environment’s responses. Each trajectory |$\ket{T}$ can be represented as |$\ket{T}=\ket{s_{0},a_{0},s^{\prime}_{0},r_{0},s_{1},a_{1},s^{\prime}_{1},r_{1},\ldots,s_{T-1},a_{T-1},s^{\prime}_{T-1},r_{T-1},g}$ where $s$ is the current state at each time step from $t=0$ to $t=T-1$, $a$ is the action taken by the agent, $s^{\prime}$ is the next state reached after taking an action at state $s$, $r$ is the reward received and $g$ is the total return throughout the trajectory.
@@ V Demonstrations and Results
In the demonstrations and results section of this work, we first present the dynamics of the classical Markov decision process (MDP) through a graphical representation of state transitions and associated probabilities, as shown in Fig. 3. The diagram includes four states labeled $s_{0},s_{1},s_{2}$ and $s_{3}$, where $s_{3}$ represents the terminal state, and two actions, $a_{0}$ and $a_{1}$. The arrows connecting the states indicate the transitions between these states for a given action. Each transition is labeled with a probability value that denotes the likelihood of moving from one state to another under a specific action. These probabilities provide insights into the stochastic nature of the state transitions. For example, transitioning from $s_{0}$ to $s_{1}$ with action $a_{0}$ has a probability of 0.6. Similarly, transitioning from $s_{0}$ to $s_{2}$ under action $a_{0}$ occurs with a probability of 0.4. The diagram also illustrates several self-transitions, such as the one in state $s_{3}$, where action $a_{1}$ leads to a transition back to $s_{3}$ with a probability of 1. These self-transitions capture situations where the system remains in the same state regardless of the action taken. The figure visualizes the probabilistic behavior of the system under different actions, as well as how the state evolves over time based on these actions.
@@ V.1.1 Agent and environment on quantum computer
Figure 4 illustrates the quantum Markov decision process (QMDP), where the agent’s actions and the environment’s responses are represented by quantum states.
@@ V.1.2 Agent and environment single interaction on quantum computer
To verify the accuracy of the QMDP circuit, we simulate a single interaction between the agent and the environment on a quantum computer. Using both a state transition heat-map and a quantum sample distribution plot, we demonstrate that the QMDP circuit correctly mirrors the state transitions and rewards of the classical MDP, mentioned in Fig. 3.
@@
Figure 6: Quantum sample distribution of the QMDP circuit, displaying the probability of measuring each quantum sample after simulating a single agent-environment interaction. The x-axis lists possible quantum samples, while the y-axis indicates their probabilities.
Figure 6 illustrates the quantum sample distribution of the QMDP circuit, displaying the probability of observing each possible quantum sample after a single execution of the circuit. Each quantum sample in binary notation (x-axis) represents a combination of the agent’s initial state, chosen action, resulting next state and reward sequence encoded in different qubits. Table 1 presents the qubit representations of each quantum sample. Starting from the least significant bits, the first two qubits represent the initial (current) state , followed by the action qubit, then the next two qubits indicate the next state, and finally, the two most significant bits represent the reward. The y-axis indicates the probability of each quantum sample. Peaks in this distribution, quantum samples |$\ket{1111111}$ and |$\ket{0101100}$ with probabilities 0.354 and 0.335 respectively, highlight high-probability sequences of this interaction. These results validate the QMDP circuit’s accurate representation of both state transition probabilities and reward during a single interaction.